# The separating Noether number of small groups

### v2

@@ Abstract.
The separating Noether number of a finite group is the minimal positive integer $d$ such that for any finite dimensional complex linear representation of the group, any two dictinct orbits can be distinguished by a polynomial invariant of degree at most $d$. In the present paper its exact value is determined for all groups of order less than $32$, for all finite groups with a cyclic subgroup of index at most $2$, and for the direct product of a dihedral group with the $2$-element group. The results show in particular that unlike the ordinary Noether number, the separating Noether number of a non-abelian finite group may well be equal to the separating Noether number of a proper direct factor of the group. Most of the results are proved for the case of a general (possibly finite) base field containing an element whose multiplicative order equals the size of the group.
@@ 1. Introduction
For a finite dimensional vector space $V$ over a field $K$ we write $K[V]$ for the coordinate ring of $V$ identified with the $n$-variable polynomial algebra $K[x_{1},\dots,x_{n}]$, where the variables $x_{1},\dots,x_{n}$ represent the coordinate functions on $V$ with respect to a chosen basis. Let $G$ be a finite group, and $\rho:G\to\mathrm{GL}(V)$ a representation of $G$ on $V$. We shall say that $(V,\rho)$ is a $G$-module in this case; mostly we shall suppress $\rho$ from the notation and we will say that $V$ is a $G$-module. Sometimes we write $V_{K}$ instead of $V$ when we want to make explicit the actual base field. The action of $G$ on $V$ induces an action on $K[V]$ via $K$-algebra automorphisms, such that the variables $x_{1},\dots,x_{n}$ span a $G$-submodule isomorphic to $V^{*}$ (the $G$-module dual to $V$), and $K[V]$ is the symmetric tensor algebra of $V^{*}$. Consider
$K[V]^{G}=\{f\in K[V]\mid\forall g\in G:\quad g\cdot f=f\},$
the subalgebra of $G$-invariants in $K[V]$. For $v\in V$ we shall denote by $G\cdot v$ the $G$-orbit of $v$, and by $\mathrm{Stab}_{G}(v):=\{g\in G\mid gv=v\}$ the stabilizer subgroup in $G$ of $v$.
It is well known that for $v,w\in V$ with $G\cdot v\neq G\cdot w$, there exists an invariant $f\in K[V]^{G}$ with $f(v)\neq f(w)$ (see for example [16, Theorem 3.12.1], [34, Theorem 16] or [23, Lemma 3.1]). Therefore the general notion of a ”separating set of invariants” (see [16, Definition 2.4.1]) reduces to the following in the case of finite groups:
Definition 1.1.
A subset $S$ of $K[V]^{G}$ is called a separating set if for any $v,w\in V$, $f(v)=f(w)$ for all $f\in S$ implies that $v$ and $w$ have the same $G$-orbit.
The result mentioned before Definition 1.1 implies that every generating set of the algebra $K[V]^{G}$ is a separating set. In the structural study of algebras of polynomial invariants of finite groups the focus on separating sets (instead of the more special generating sets) turned out to be useful in finding strengthened versions of known theorems in the invariant theory of finite groups, see for example [24], [25]. Considering explicit computations of invariants, the class of modular representations for which generators of the corresponding algebra of invariants is known is very limited. On the other hand, in several such cases explicit finite separating sets are known, see for example [49], [38]. The minimal possible size of a separating set is investigated in [26]. In the non-modular case, an explicit separating set of multisymmetric polynomials is given in [40]. The systematic study of separating invariants over finite fields began recently in [35], [39].
1.1. Degree bounds
Endow $K[V]$ with the standard grading, so the variables $x_{1},\dots,x_{n}$ span the degree $1$ homogeneous component of $K[V]$. We shall denote by $\deg(f)$ the degree of a non-zero homogeneous element of $K[V]$. Obviously, $K[V]^{G}$ is spanned by homogeneous elements, so it inherits the grading from $K[V]=\bigoplus_{d=0}^{\infty}K[V]_{d}$. We shall write $K[V]^{G}_{\leq d}$ for the sum of the homogeneous components of degree at most $d$. Degree bounds for generators of algebras of invariants have always been in the focus of invariant theory (see e.g. [50]). We set
$\beta(G,V):=\min\{d\in\mathbb{N}_{0}\mid K[V]^{G}_{\leq d}\text{ generates the}K\text{-algebra }K[V]^{G}\},$
$\beta^{K}(G):=\sup\{\beta(G,V)\mid V\text{ is a finite dimensional }G\text{-module}\text{ over }K\}.$
The quantity $\beta^{K}(G)$ is called the Noether number of $G$ (over $K$) because of the classical inequality $\beta^{\mathbb{C}}(G)\leq|G|$ due to E. Noether [41] (the inequality was extended to the non-modular case in [29], [30]). Its systematic study was initiated in [47]. The results of [47] were improved first in [19], [48] and later in [11], [12]. In particular, [15, Table 1] gives the exact value of the Noether number for all non-abelian groups of order less than $32$, and [12] gives the Noether number for all groups containing a cyclic subgroup of index two.
Following [37, Definition 2] (see also [34]) we set
$\beta_{\mathrm{sep}}(G,V):=\min\{d\in\mathbb{N}_{0}\mid K[V]^{G}_{\leq d}\text{ is a separating set}\},$
and
$\beta_{\mathrm{sep}}^{K}(G):=\sup\{\beta_{\mathrm{sep}}(G,V)\mid V\text{ is a finite dimensional }G\text{-module}\text{ over }K\}.$
We shall refer to the above quantities as the separating Noether number. As different orbits of a finite group can be separated by polynomial invariants, we have the obvious inequality
$\beta_{\mathrm{sep}}(G,V)\leq\beta(G,V)\qquad\text{ for any }V,$
and hence
(1) $\beta_{\mathrm{sep}}^{K}(G)\leq\beta^{K}(G).$
A major difference between the Noether number and its separating version is that if the characteristic of the base field $K$ divides the order of $G$, then $\beta^{K}(G)=\infty$ by [43], whereas $\beta_{\mathrm{sep}}^{K}(G)\leq|G|$ also in the modular case (see for example the proof of [16, Theorem 3.12.1], [34, Theorem 16] or [23, Lemma 3.1]) However, in the present paper our focus is on the non-modular setup, when the characteristic of the base field does not divide the order of $G$. In this context an interesting and generally open question is the following:
Question 1.2.
To what extent is the change from ”generating sets” to ”separating sets” reflected in the corresponding degree bounds?
Moreover, interest in these degree bounds arose also in recent applications for data science and machine learning, see [2], [5], [6], [7]. We mention also that problems in signal processing motivated the study of a variant of the Noether number for rational invariants in [4].
The exact value of $\beta_{\mathrm{sep}}^{K}(G)$ for certain finite abelian groups $G$ and an algebraically closed base field of non-modular characteristic was determined recently in [45], [46]. As far as we know, the only non-abelian finite groups whose separating Noether number in the non-modular case is determined in the literature are the non-abelian groups of order $3p$ for some prime $p$ congruent to $1$ modulo $3$ (see [9]). For some results on the modular case see [37, Theorem C], [27].
The aim of the present work is to produce a counterpart for the separating Noether number of the results on the Noether number from [12] and [15].
Running assumption. Throughout this paper $G$ will stand for a finite group, and $K$ for a field. Unless explicitly stated otherwise, it is assumed that the characteristic of $K$ does not divide $|G|$.
2. Main results
Our first result builds on [12] (see the equality (13)) and shows that for a group with a cyclic subgroup of index $2$ (their classification will be recalled in Section 4) the inequality (1) holds with equality, namely:
@@ 2. Main results
Let $G$ be a non-cyclic finite group with a cyclic subgroup of index $2$, and assume that $K$ contains an element of multiplicative order $|G|$. Then we have the equality
$\beta_{\mathrm{sep}}^{K}(G)=\frac{1}{2}|G|+\begin{cases}2&\text{ if }G=\mathrm{Dic}_{4m},\text{ $m>1$};\\ 1&\text{ otherwise. }\end{cases}$
Using [12, Corollary 5.5] we are able to determine the separating Noether number for another infinite sequence of non-abelian finite groups:
Theorem 2.2.
For $n\geq 2$ even consider the direct product $G:=\mathrm{D}_{2n}\times\mathrm{C}_{2}$ of the dihedral group of order $2n$ and the cyclic group $\mathrm{C}_{2}$ of order $2$. Assume that $K$ has an element with multiplicative order $n$. Then we have
$\beta_{\mathrm{sep}}^{K}(\mathrm{D}_{2n}\times\mathrm{C}_{2})=n+2.$
There are $16$ non-abelian groups of order less than $32$ not covered by Theorem 2.1 or Theorem 2.2. We discuss them in the following theorem:
Theorem 2.3.
Let $G$ be a non-abelian group of order less than $32$ that does not contain a cyclic subgroup of index at most $2$, and $G$ is not isomorphic to $\mathrm{D}_{2n}\times\mathrm{C}_{2}$. Assume that $K$ contains an element of multiplicative order $|G|$. The value of $\beta_{\mathrm{sep}}^{K}(G)$ is given in the following table, in some cases under the additional assumption that $\mathrm{char}(K)=0$ (this is indicated in the corresponding line of the table):
@@
The first column in the above table contains the identifier of $G$ in the Small Groups Library of GAP (see [31]); that is, a pair of the form [order, i], where the GAP command $\mathsf{SmallGroup(id)}$ returns the i-th group of order order in the catalogue. The second column gives $G$ in a standard notation or indicates its structure as a direct product or semidirect product (the symbol $\rtimes$ always stands for a semidirect product that is not a direct product, and $\psi$ and $\phi$ in the rows $(16,3)$ and $(16,13)$ stand for two different involutive automorphisms of $\mathrm{C}_{4}\times\mathrm{C}_{2}$). The presentation of $G$ in terms of generators and relations can be found in the beginning of the section that contains the theorem referred to in the last column for the value of $\beta_{\mathrm{sep}}^{K}(G)$. The third column of the table contains $\beta^{K}(G)$; these values are taken from [15]. The fourth column contains $\beta_{\mathrm{sep}}^{K}(G)$.
Note that in some cases the value for $\beta_{\mathrm{sep}}^{K}(G)$ given in the table holds under weaker restrictions on $K$ than the restrictions stated in Theorem 2.3. The exact conditions on $K$ needed for the validity of our arguments are contained in the statement quoted in the last column of the table. In a few instances we used computer calculations to determine generators of concrete invariant algebras. This reduces the validity of some of our results to the case of a base field of characteristic zero. The use of computer could have been avoided at the price of a significant increase of the length of the paper.
For sake of completeness we include a similar table containing $\beta_{\mathrm{sep}}^{K}(G)$ for those abelian groups $G$ of order less than $32$ that do not have a cyclic subgroup of index at most $2$, under the assumption that $K$ contains an element whose multiplicative order equals the exponent of $G$. These results follow from [18], [45], [46], and the explanations in Section 3.1.
[…]
2.1. Failure of strict monotonicity
For a normal subgroup $N$ of $G$, all representations of $G/N$ can be viewed as a representation of $G$ whose kernel contains $N$. Therefore we have the obvious inequalities
(2) $\beta^{K}(G)\geq\beta^{K}(G/N)\text{ and }\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(G/N).$
Moreover, both the Noether number and the separating Noether number are monotone for subgroups as well; that is, for any subgroup $H$ of a finite group $G$ we have
(3) $\beta^{K}(G)\geq\beta^{K}(H)$
(see [47]) and
(4) $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(H)$
(see [37, Theorem B]). The inequality (3) was sharpened in [13, Theorem 1.2], where the following strict monotonicity of the Noether number was proved:
(5) $\beta^{K}(G)>\beta^{K}(H)\text{ for any subgroup }H\neq G;$
$\beta^{K}(G)>\beta^{K}(G/N)\text{ for any normal subgroup }N\neq\{1_{G}\}\text{ of }G.$
The results of Theorem 2.1 and Theorem 2.3 show that the analogues of (5) (i.e. strict monotonicity) does not hold for the separating Noether number. Indeed, for $K$ as in the cited theorems, we have
$\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8})=6=\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8}\times\mathrm{C}_{2}),$
and $\mathrm{Dic}_{8}\times\mathrm{C}_{2}$ has a proper direct factor isomorphic to $\mathrm{Dic}_{8}$. Another similar example is
$\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4})=6=\beta_{\mathrm{sep}}^{K}(\mathrm{A}_{4}\times\mathrm{C}_{2}).$
Furthermore,
$\beta_{\mathrm{sep}}^{K}(\mathrm{S}_{3}\times\mathrm{C}_{3})=6=\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{6}),$
and the group $\mathrm{S}_{3}\times\mathrm{C}_{3}$ has a subgroup isomorphic to $\mathrm{C}_{6}$. Similar examples exist among the abelian groups. For example, we have
$\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{6}\oplus\mathrm{C}_{6}\oplus\mathrm{C}_{6})=12=\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{6}\oplus\mathrm{C}_{6}\oplus\mathrm{C}_{3})$
by [45, Theorem 6.1].
2.2. Comments on the methods used
@@
We make an essential use of the known results on exact values of the Noether numbers (as summarized in [15]) of the groups considered here. In particular, in each case when the Noether number and the separating Noether number happen to coincide, we just need to come up with a representation and two points with distinct orbits that can not be separated by invariants of strictly smaller degree than the Noether number. In all such cases it turns out that rather small dimensional representations suffice to provide such examples. On the other hand, to settle the cases when the separating Noether number is strictly smaller than the Noether number requires elaborate work: we need a thorough analysis of concrete representations, discussions of stabilizers, group automorphisms, construction of invariants and relative invariants, as well as several ad hoc ideas that help to undertand the set of solutions of systems of polynomial equations. What makes these computations feasible is that by general principles (see the notion of Helly dimension and Lemma 3.7 below) it is sufficient to deal with representations having a small number of irreducible summands.
2.3. Organization of the paper
We begin in Sections 3.1 and 3.2 with a discussion of the dependence of $\beta_{\mathrm{sep}}^{K}(G)$ on the base field $K$, which seems to be stronger than the dependence of $\beta^{K}(G)$ on $K$. In the rest of Section 3 we introduce notation and collect general facts used throughout the paper. We turn in Section 4 to the proof of Theorem 2.1 and Theorem 2.2. In Section 5 we deal with the groups of order less than $32$ for which the separating Noether number equals the Noether number. In Section 6 we cover three groups of the form $H\times\mathrm{C}_{2}$ satisfying $\beta_{\mathrm{sep}}^{K}(H\times\mathrm{C}_{2})=\beta_{\mathrm{sep}}^{K}(H)$. Section 7 is devoted to the study of the separating Noether number of $\mathrm{S}_{3}\times\mathrm{C}_{3}$. In the final Section 8 the separating Noether number is computed for four groups of the form $G\cong\mathrm{C}_{n}\rtimes\mathrm{C}_{k}$ with $\beta_{\mathrm{sep}}^{K}(G)<\beta^{K}(G)$.
@@ 3. Preliminaries
3.1. Dependence on the base field of $\beta_{\mathrm{sep}}^{K}(G)$

### v3

@@ Abstract.
The present paper completes the computation of the separating Noether numbers for the groups with order strictly less than $32$. Most of the results are proved for the case of a general (possibly finite) base field containing an element whose multiplicative order equals the size of the group.
@@
Partially supported by the Hungarian National Research, Development and Innovation Office, NKFIH K 138828.
@@ 1. Introduction
Throughout this paper $G$ is a finite group, and $K$ is a field whose characteristic does not divide $|G|$. Given a linear action of $G$ on an $n$-dimensional $K$-vector space $V$ (i.e. $V$ is a $KG$-module), denote by $K[V]^{G}$ the subalgebra of the polynomial ring $K[V]=K[x_{1},\dots,x_{n}]$ consisting of $G$-invariant polynomials. Since the action of $G$ on $K[V]$ preserves the standard grading, the algebra $K[V]^{G}$ is generated by homogeneous elements. Write $\beta(G,V)$ for the minimal positive integer $d$ such that $K[V]^{G}$ is generated by elements of degree at most $d$, and denote by $\beta^{K}(G)$ the supremum of the numbers $\beta(G,V)$, as $V$ ranges over all finite dimensional $KG$-modules. A classical theorem of Noether [28] asserts that $\beta^{\mathbb{C}}(G)\leq|G|$, and therefore $\beta^{K}(G)$ is called the Noether number of $G$ (over $K$). The systematic study of $\beta^{K}(G)$ was initiated in [34], and continued in [14], [35], [7], [8], [5]. As far as the exact value of $\beta^{K}(G)$ is concerned, the state of the art is summarized in [11]. Note that for an abelian group $G$, the Noether number $\beta^{\mathbb{C}}(G)$ equals the Davenport constant of $G$, which has an extensive literature, see for example [36] for a recent work.
Here we investigate a similar quantity, called the separating Noether number. It is related to one of the main motivations for studying $K[V]^{G}$, namely to the desire to distinguish the $G$-orbits in $V$. Denote by $G\cdot v$ the $G$-orbit of $v\in V$. For $v,w\in V$ with $G\cdot v\neq G\cdot w$, there exists an invariant $f\in K[V]^{G}$ with $f(v)\neq f(w)$ (see for example [12, Theorem 3.12.1], [23, Theorem 16] or [16, Lemma 3.1]). Therefore we call a subset $S$ of $K[V]^{G}$ a separating set of $G$-invariants (separating set for short) if whenever $G\cdot v\neq G\cdot w$, then there exists an $f\in S$ with $f(v)\neq f(w)$.
The separating Noether number $\beta_{\mathrm{sep}}(G,V)$ is the minimal positive integer $d$ such that the elements of degree at most $d$ in $K[V]$ form a separating set, and $\beta_{\mathrm{sep}}^{K}(G)$ is the supremum of $\beta_{\mathrm{sep}}(G,V)$, as $V$ ranges over all finite dimensional $KG$-modules (these quantities were defined in [25], [23]). Clearly, any generating set is a separating set, hence we have the obvious inequalitites
(1) $\beta_{\mathrm{sep}}(G,V)\leq\beta(G,V)\text{ and }\beta_{\mathrm{sep}}^{K}(G)\leq\beta^{K}(G).$
The study of separating sets of polynomial invariants became a popular topic in the past decades, see [17], [18], [23], [16], [27], [25], [26], [24] for a sample of papers on this subject. A recent interest in separating invariants arose in machine learning and signal processing, see for example [1], [3], [4], [19], [2].
In [15] we initiated the following program: determine the separating Noether number for all the non-abelian groups whose Noether number is known (i.e. the groups covered by [11]). The present paper completes this task. The case of abelian groups is covered by [13], [31], [32], [33]. Interestingly, the range of finite abelian groups whose separating Noether number is known goes far beyond the range of finite abelian groups with known Noether number (Davenport constant). The non-abelian groups of order at most $16$, as well as the groups with a cyclic subgroup of index $2$ and the direct products of dihedral groups with the $2$-element group are dealt with in [15]. It remains to handle the non-abelian groups $G$ with $16<|G|<32$ which do not fit into the infinite series discussed in [15]. This is the content of the present paper.
1.1. The main result
@@
Let $G$ be a non-cyclic group with $16<|G|<32$, and assume that $K$ contains an element of multiplicative order $|G|$. The value of $\beta_{\mathrm{sep}}^{K}(G)$ is given in Table 1, in some cases under the additional assumption that $\mathrm{char}(K)=0$ (this is indicated in the corresponding line of the table).
Table 1. Separating Noether numbers for $16<|G|<32$
@@
The new parts of Theorem 1.1 are the values of $\beta_{\mathrm{sep}}(G)$ for the $10$ non-abelian groups in Table 1 that do not contain a cyclic subgroup of index $2$, and are not isomorphic to $\mathrm{D}_{2n}\times\mathrm{C}_{2}$ or to $\mathrm{C}_{7}\rtimes\mathrm{C}_{3}$. The first column of Table 1 contains the identifier of $G$ in the Small Groups Library of GAP (see [20]); that is, a pair of the form (order, i). The GAP command $\mathsf{SmallGroup(id)}$ returns the i-th group of order order in the catalogue. The second column gives $G$ in a standard notation or indicates its structure as a direct product or semidirect product. The symbol $\rtimes$ always stands for a semidirect product that is not a direct product ($\gamma$ in the row $(24,8)$ stands for an involutive automorphisms of $\mathrm{C}_{6}\times\mathrm{C}_{2}$ different from $g\mapsto g^{-1}$). The third column of the table is filled in with the values of $\beta^{K}(G)$ (taken from [11]). The fourth column contains $\beta_{\mathrm{sep}}^{K}(G)$. In the last column one can find the references for the theorems discussing the value of $\beta_{\mathrm{sep}}^{K}(G)$.
1.2. Comments on the base field, on computer usage
By Lemma 2.3 below, the value in the column of $\beta_{\mathrm{sep}}^{K}(G)$ in Table 1 is an upper bound for the separating Noether number under the assumption that $\mathrm{char}(K)$ does not divide |G||G|” (which is weaker than the assumption that ”$K$ contains an element of multiplicative order |G||G|” in Theorem 1.1).
Standard computer algebra packages can compute a minimal system of generators for rings of invariants of finite groups. We made a limited use of the online CoCalc platform [30] for this purpose. Namely, we verified by it some desired upper bounds for $\beta(G,V)$ for certain representations of the groups $\mathrm{C}_{5}\rtimes\mathrm{C}_{4}$, $\mathrm{Dic}_{12}\times\mathrm{C}_{2}$, and $\mathrm{M}_{27}$, in order to avoid a significant increase in the length of the paper. In these cases we did the calculations only in the most classical case of a characteristic zero base field (we could not settle all the infinitely many non-modular characteristics by computer anyway), that is why for these groups we need to assume $\mathrm{char}(K)=0$.
We are not aware of a computer algebra package that computes the separating Noether number for a finite group. A simple algorithm for algebraically closed base fields is sketched in [15], and we used an implementation of it on [30] to verify over the complex base field the results of [15] on the groups of order at most $16$. The same computational power turned out to be insufficient to deal with the groups studied in the present paper. Moreover, the results of the present paper in full power (valid in all but finitely many characteristics, for infinitely many different fields that are not necessarily algebraically closed) could not be obtained by computer calculations, regardless of the available computational power.
@@ 1.2. Comments on the base field, on computer usage
In Section 2 we introduce notation and mention general facts used throughout the paper. In Section 3 we deal with those groups from Theorem 1.1 for which the separating Noether number equals the Noether number. By the inequality (1) in these cases it suffices to give a representation and two points with distinct orbits that can not be separated by invariants of strictly smaller degree than the Noether number. The cases when the separating Noether number is strictly smaller than the Noether number require more elaborate work, and this is done in the remaining Sections: we need a thorough analysis of concrete representations, discussions of stabilizers, group automorphisms, construction of invariants and relative invariants, as well as several ad hoc ideas that help to understand the set of solutions of systems of polynomial equations. What makes these computations feasible is that general principles imply that it is sufficient to deal with representations having a small number of irreducible direct summands (see [15, Lemma 6.2] about the Helly dimension).
@@ 2. Preliminaries
2.1. Notational conventions.
In the sequel we shall typically define an irreducible $KG$-module $V$ for a group $G$ given by generators and relations as follows. We specify a group homomorphism $\psi:G\to\mathrm{GL}_{n}(K)$ by giving the images of the generators. Then $V=K^{n}$ is the space of column vectors, and $g\cdot v:=\psi(g)v$ (matrix multiplication on the right hand side). The induced action of $G$ on the variables $x_{1},\dots,x_{n}$ (which is a basis in the dual space $V^{*}$ of $V$) is given by
$g\cdot x_{j}=\sum_{i=1}^{n}\psi(g^{-1})_{ji}x_{i}.$
Let $\widehat{G}=\mathrm{Hom}(G,K^{\times})$ be the character group of $G$. We denote by $U_{\chi}$ the $1$-dimensional vector space corresponding to the element $\chi\in\widehat{G}$; that is, $g\cdot\lambda=\chi(g)\lambda$ for $g\in G$ and $\lambda\in K=U_{\chi}$. Denote by $t_{\chi}$ the coordinate function on $U_{\chi}$. Therefore we have $g\cdot t_{\chi}=\chi(g^{-1})t_{\chi}$.
Usually we shall consider a $KG$-module $V$ with a direct sum decomposition
where the $W_{i}$ are pairwise non-isomorphic irreducible $KG$-modules of dimension at least $2$, and $\chi_{1},\dots,\chi_{m}$ are distinct characters of $G$. The coordinate functions on $W_{1},W_{2}$ respectively $W_{3}$, etc, will be denoted by $x_{1},x_{2},\dots$, $y_{1},y_{2},\dots$, respectively $z_{1},z_{2},\dots$, etc.
Associated to the above direct sum decomposition (2) of $V$ there is an $\mathbb{N}_{0}^{l+m}$-grading on $K[V]$. Namely, the component of multidegree $\alpha=(\alpha_{1},\dots,\alpha_{l+m})$ of $K[V]$ consists of the polynomials homogeneous of degree $\alpha_{j}$ on the $j$th direct summand of $V=W_{1}\oplus\cdots\oplus U_{\chi_{m}}$ for each $j=1,\dots,l+m$. This multigrading is induced by the following action of the torus $T:=K^{\times}\times\cdots\times K^{\times}$ ($l+m$ direct factors) on $V$: for $\lambda=(\lambda_{1},\dots,\lambda_{l+m})\in T$ and $v=(w_{1},\dots,w_{l},u_{1},\dots,u_{m})\in V$ set $\lambda\cdot v:=(\lambda_{1}w_{1},\dots,\lambda_{l+m}u_{m})$. This action commutes with the $G$-action on $V$, and therefore $K[V]^{G}$ is spanned by multihomogeneous elements. Moreover, for $v,v^{\prime}\in V$ we have that
When we say in the text that ’after rescaling, we may assume that $v$ has some special form’, then we mean that we replace $v$ by an appropriate element in its $T$-orbit, and we refer to (3). The phrase ‘multihomogeneous’ will refer to this multigrading in the text.
2.2. Relative invariants
For a $KG$-module $V$, denote by $K[V]^{G}_{+}$ the sum of the positive degree homogeneous components of $K[V]^{G}$. A homogeneous invariant $f\in K[V]^{G}$ is indecomposable if $f\notin(K[V]^{G}_{+})^{2}$. Clearly, $K[V]^{G}$ is generated as an algebra by the positive degree homogeneous indecomposable invariants. The Hilbert ideal $\mathcal{H}(G,V)$ is the ideal in the polynomial algebra $K[V]$ generated by the positive degree homogeneous $G$-invariants.
We say that an element $f\in K[V]$ is a relative invariant of weight $\chi\in\widehat{G}$ if $g\cdot f=\chi(g^{-1})f$ for all $g\in G$ and set $K[V]^{G,\chi}:=\{f\in K[V]\mid f\text{ is a relative invariant of weight }\chi\}$. For example, the coordinate function $t_{\chi}$ on $U_{\chi}$ is a relative invariant of weight $\chi$. Observe that for a sequence $\chi^{(1)},\dots,\chi^{(k)}$ of elements of $\widehat{G}$, $t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}$ is an indecomposable $G$-invariant if and only if $\chi^{(1)},\dots,\chi^{(k)}$ is an irreducible product-one sequence over $\widehat{G}$, i.e. $\chi^{(1)}\cdots\chi^{(k)}=1\in\widehat{G}$ and there is no $\ell<k$, $1\leq i_{1}<\cdots<i_{\ell}\leq k$ with $\chi^{(i_{1})}\cdots\chi^{(i_{\ell})}=1\in\widehat{G}$. Here $k$ is the length of the sequence $\chi^{(1)},\cdots,\chi^{(k)}$. We say that $\chi^{(1)},\dots,\chi^{(k)}$ is a product-one free sequence over $\widehat{G}$ if there is no $1\leq i_{1}<\dots<i_{\ell}\leq k$ with $\chi^{(i_{1})}\cdots\chi^{(i_{\ell})}=1\in\widehat{G}$. The maximal length of an irreducible product-one sequence over $G$ is called the Davenport constant $\mathsf{D}(G)$ of the group $G$. For a finite abelian group $G$, the maximal possible length of a product-one free sequence is $\mathsf{D}(G)-1$. For more information about the Davenport constant and its relevance for the invariant theory of abelian groups see for example [10].
In some cases we will compute a homogeneous generating system of $K[V]^{G}$ by using the following observation: